OpenStudy (anonymous):

The maximum speed of a coyote is 43 miles per hours. What is a coyote's maximum speed in feet per second?Round answer to the nesrest tenth.(1 mile=5280ft;1hr=3600 sec)

OpenStudy (anonymous):

\[\frac{43*5280}{60*60}=\frac{946}{15}=63.07 \text{ ft}/\sec \]

OpenStudy (anonymous):

60 mph = 88 ft/sec 43/60 * 88 = 946/15 = 63.07 ft/sec